# On novel Hamiltonian descriptions of some three-dimensional non-conservative systems

### v1

@@
Aritra Ghosh [^1] Corresponding author: aritraghosh500@gmail.com, ag34@iitbbs.ac.in, Anindya Ghose-Choudhury222aghosechoudhury@gmail.com, Partha Guha333partha.guha@ku.ac.ae
@@ 2 Hamiltonian dynamics on Poisson manifolds: A review
$[\Lambda,\Lambda]=0,$ (2.1)
with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] being the Schouten-Nijenhuis bracket [2]. For two functions $f,g\in C^{\infty}(\mathcal{M},\mathbb{R})$, the Poisson bi-vector allows one to define a local Lie bracket on $C^{\infty}(\mathcal{M},\mathbb{R})$ as $\Lambda(df,dg)=\{f,g\}$ called the Poisson bracket which is obviously skew-symmetric. Additionally, the condition (2.1) enforces the so-called Jacobi identity $\{\{f,g\},h\}+\{\{h,f\},g\}+\{\{g,h\},f\}=0$, $\forall f,g,h\in C^{\infty}(\mathcal{M},\mathbb{R})$.
@@
where $\mathbf{x}=(x_{1}\quad x_{2}\cdots x_{n})^{T}$ and $\nabla$ is the usual del operator. An immediate consequence of (2.2) or (2.5) is that $\frac{dH}{dt}=X_{H}(H)=0$, i.e., the Hamiltonian $H$ is conserved. Thus, Hamiltonian dynamics on Poisson manifolds describes conservative systems that conserves the Hamiltonian function; notice that symplectic manifolds encountered in classical mechanics are all endowed with a Poisson structure due to the symplectic two-form [1]. It is then naturally implied that the description of non-conservative systems is possible only upon introducing additional elements into the formalism.
@@ 2.1 Three-dimensional systems
Notice that the general solution of the equation (2.7) is of the form (see [21] and references therein)
@@
for some function $\overline{H}\in C^{\infty}(\mathcal{M},\mathbb{R})$ and $M$ is an integrating factor called the Jacobi last multiplier [28, 29]. Thus, a Poisson system may admit a non-trivial divergence [^4] For an arbitrary vector field $X$ on a smooth and orientable manifold $\mathcal{M}$ with volume-form $\Omega=dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{n}$, the divergence can be defined as £$\pounds_{X}\Omega={\rm div}(X)\Omega$ (see [36] and references therein). In the notation of three-dimensional vectors in $\mathbb{R}^{3}$ as employed in this paper, this agrees with that obtained by $\nabla\cdot\dot{\mathbf{x}}$. unlike a Hamiltonian system on a symplectic manifold [1] because
@@
i.e., if $M$ is not a constant, then $\nabla\cdot\dot{\mathbf{x}}\neq 0$ and vice versa; notice that $M$ is a constant if and only if ×$\nabla\times\mathbf{J}=0$. The above discussion reveals that we can re-write the vector equation (2.8) using (2.9) such that
@@
[…]
@@ 3 Resistive and port-Hamiltonian systems
where we have used Ohm’s law to write the voltage drop across the resistance $R$ as $-IR$. The system (3.3) can no longer be expressed in the form of (2.5) for a skew-symmetric matrix $\mathcal{J}$; however, one can write down a modified evolution equation for the same Hamiltonian $H=\frac{I^{2}}{2}+\frac{Q^{2}}{2LC}$ as
@@ 3.1.1 Reduced three-wave-interaction model
The reduced three-wave-interaction model arises in physical situations where three quasisynchronous waves interact in a plasma with quadratic nonlinearities. The system is described by the following set of first-order equations:
@@
the last result indicating that the dynamics does not preserve the volume-form $\Omega=dx\wedge dy\wedge dz$ in $\mathbb{R}^{3}$. One can check that the corresponding Poisson vector is irrotational, i.e., ×$\nabla\times\mathbf{J}=0$, meaning that the divergence above is purely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.1.2 Rabinovich system
meaning that the volume-form $\Omega=dx\wedge dy\wedge dz$ in $\mathbb{R}^{3}$ is non-conserved. It is easy to check that the corresponding Poisson vector is irrotational, i.e., ×$\nabla\times\mathbf{J}=0$, implying that the divergence above is purely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.2.1 Chen system
One can verify that ×$\nabla\times\mathbf{J}=0$, meaning that the Jacobi identity (2.7) is automatically satisfied but this moreover indicates that the non-trivial divergence of the dynamical vector field $\mathcal{X}_{H}$ is entirely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.2.2 Lü system
It may be checked that ×$\nabla\times\mathbf{J}=0$, implying that the Jacobi identity (2.7) is automatically satisfied but this moreover indicates that the non-trivial divergence of the dynamical vector field $\mathcal{X}_{H}$ is entirely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.2.3 Modified Lü system
In this case, one has ×$\nabla\times\mathbf{J}\neq 0$ but $\mathbf{J}\cdot(\nabla\times\mathbf{J})=0$, meaning that the Jacobi identity (2.7) is satisfied.
@@ 3.2.4 Qi system
In this case one has ×$\nabla\times\mathbf{J}\neq 0$ but $\mathbf{J}\cdot(\nabla\times\mathbf{J})=0$, meaning that the Jacobi identity (2.7) is satisfied.
@@ 4.1.2 Lü system
Although the Poisson vector $\mathbf{N}=(N_{x},N_{y},N_{z})$ satisfies the condition (×$\nabla\cdot(\nabla\times\mathbf{N})=0$ for arbitrary values of the underlying parameters, let us pick $\beta=\gamma$ for the sake of simplicity as this gives ×$\nabla\times\mathbf{N}=0$ and consequently from (4.5) we get $M=1$ and
@@ 4.2 Jordan-like structures
This opens up certain possibilities; most importantly that of generating new bi-Hamiltonian systems guided by the structure of Jordan homomorphisms [^6] Recall that a map between two rings $\phi:R\rightarrow R^{\prime}$ is a Jordan homomorphism if (a) $\phi(a+b)=\phi(a)+\phi(b)$, (b) $\phi(ab+ba)=\phi(a)\phi(b)+\phi(b)\phi(a)$, for all $a,b\in R$. In the present case, although we do not get a Jordan algebra, we can make use of transformations inspired from Jordan homomorphisms which we will call Jordan-like transformations. Let us consider an invertible matrix $\mathcal{T}$ under which $\mathcal{J}$ and $\mathcal{R}$ transform as
@@
Note that upon putting $\Delta=0$, one recovers the bi-Hamiltonian system (4.11). The system admits a non-trivial Jacobi last multiplier which may be evaluated straightforwardly and this is therefore not done here. In summary, Jordan-like transformations that map $\mathcal{N}=\mathcal{J}\mathcal{R}+\mathcal{R}\mathcal{J}$ to another Poisson matrix $\mathcal{N}^{\prime}$ can be used to generate new bi-Hamiltonian systems as illustrated by the simple example above.
@@ 5 Conformal Hamiltonian dynamics on Poisson manifolds
£$\pounds_{X^{a}_{H}}\Omega=(na)\Omega,$ (5.3)
@@
Note that upon putting $a=0$ one recovers the standard Hamiltonian vector field. The Hamiltonian is not conserved under the flow of $X^{a}_{H}$ for $a\neq 0$ as $X^{a}_{H}(H)=a\Gamma(H)$. Thus, a conformal Hamiltonian vector field has the typical characteristics required for describing a non-conservative time evolution. A key feature of a conformal Hamiltonian vector field is the constancy of its divergence, i.e., although the divergence is non-zero, it is a constant. In what follows, we will demonstrate how some of our previously-considered examples can be described alternatively using conformal Hamiltonian dynamics on Poisson manifolds, i.e., without restoring to resistance matrices.
@@ 5.1 Reduced three-wave-interaction model
The phase trajectories of the system (3.9) are then the integral curves of the conformal Hamiltonian vector field $X^{-2}_{H}$, i.e., with $a=-2$ in (5.2).
@@ Appendix A Reverse-engineering non-conservative systems
Obviously, the Jacobi identity is satisfied. The Euler equations are obtained from $\dot{\mathbf{x}}=\mathcal{N}(\nabla G)$, where $\mathbf{x}=(L_{x}\quad L_{x}\quad L_{z})^{T}$ and $G$ is the identified with the total kinetic energy $K$. The result is the familiar system of equations

### v2

@@
Aritra Ghosh111aritraghosh500@gmail.com, ag34@iitbbs.ac.in, Anindya Ghose-Choudhury222aghosechoudhury@gmail.com, Partha Guha333partha.guha@ku.ac.ae
@@ 2 Hamiltonian dynamics on Poisson manifolds: A review
$\Lambda,\Lambda]=0,$ (2.1)
with $\cdot$$\cdot$] being the Schouten-Nijenhuis bracket [2]. For two functions $f,g\in C^{\infty}(\mathcal{M},\mathbb{R})$, the Poisson bi-vector allows one to define a local Lie bracket on $C^{\infty}(\mathcal{M},\mathbb{R})$ as $\Lambda(df,dg)=\{f,g\}$ called the Poisson bracket which is obviously skew-symmetric. Additionally, the condition (2.1) enforces the so-called Jacobi identity $\{\{f,g\},h\}+\{\{h,f\},g\}+\{\{g,h\},f\}=0$, $\forall f,g,h\in C^{\infty}(\mathcal{M},\mathbb{R})$.
@@
where $\mathbf{x}=(x_{1}\quad x_{2}\cdots x_{n})^{T}$ and $\nabla$ is the usual del operator. An immediate consequence of (2.2) or (2.5) is that $\frac{dH}{dt}=X_{H}(H)=0$, i.e., the Hamiltonian $H$ is conserved. Thus, Hamiltonian dynamics on Poisson manifolds describes conservative systems that conserve the Hamiltonian function; notice that symplectic manifolds encountered in classical mechanics are all endowed with a Poisson structure due to the symplectic two-form [1]. It is then naturally implied that the description of non-conservative systems is possible only upon introducing additional elements into the formalism.
@@ 2.1 Three-dimensional systems
Notice that the general solution of the equation (2.7) is locally of the form (see [21] and references therein)
@@
for some function $\overline{H}\in C^{\infty}(\mathcal{M},\mathbb{R})$ and $M$ is an integrating factor called the Jacobi last multiplier [28, 29]. Thus, a Poisson system may admit a non-trivial divergence [^4] For an arbitrary vector field $X$ on a smooth and orientable manifold $\mathcal{M}$ with volume-form $\Omega=dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{n}$, the divergence can be defined as $\pounds_{X}\Omega={\rm div}(X)\Omega$ (see [36] and references therein). In the notation of three-dimensional vectors in $\mathbb{R}^{3}$ as employed in this paper, this agrees with that obtained by $\nabla\cdot\dot{\mathbf{x}}$. unlike a Hamiltonian system on a symplectic manifold [1] because
@@
i.e., if $M$ is not a constant, then $\nabla\cdot\dot{\mathbf{x}}\neq 0$ and vice versa; notice that $M$ is a constant if and only if $\nabla\times\mathbf{J}=0$. The above discussion reveals that we can re-write the vector equation (2.8) using (2.9) such that
@@
where $\{\cdot,\cdot,\cdot\}$ is the Nambu three-bracket which satisfies a generalized and stronger version of the Jacobi identity called the fundamental identity which was introduced in [12]. Note that any three-dimensional bi-Hamiltonian system is also a Nambu-Poisson system.
@@ 3 Resistive and port-Hamiltonian systems
where we have used Ohm’s law to write the voltage drop across the resistance $R$ as IR-IR. The system (3.3) can no longer be expressed in the form of (2.5) for a skew-symmetric matrix $\mathcal{J}$; however, one can write down a modified evolution equation for the same Hamiltonian $H=\frac{I^{2}}{2}+\frac{Q^{2}}{2LC}$ as
@@ 3.1.1 Reduced three-wave-interaction model
The reduced three-wave-interaction model arises in physical situations where three quasi-synchronous waves interact in a plasma with quadratic nonlinearities. The system is described by the following set of first-order equations:
@@
the last result indicating that the dynamics does not preserve the volume-form $\Omega=dx\wedge dy\wedge dz$ in $\mathbb{R}^{3}$. One can check that the corresponding Poisson vector is irrotational, i.e., $\nabla\times\mathbf{J}=0$, meaning that the divergence above is purely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.1.2 Rabinovich system
meaning that the volume-form $\Omega=dx\wedge dy\wedge dz$ in $\mathbb{R}^{3}$ is non-conserved. It is easy to check that the corresponding Poisson vector is irrotational, i.e., $\nabla\times\mathbf{J}=0$, implying that the divergence above is purely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.2.1 Chen system
One can verify that $\nabla\times\mathbf{J}=0$, meaning that the Jacobi identity (2.7) is automatically satisfied but this moreover indicates that the non-trivial divergence of the dynamical vector field $\mathcal{X}_{H}$ is entirely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.2.2 Lü system
It may be checked that $\nabla\times\mathbf{J}=0$, implying that the Jacobi identity (2.7) is automatically satisfied but this moreover indicates that the non-trivial divergence of the dynamical vector field $\mathcal{X}_{H}$ is entirely due to the resistive part of the time evolution as described by the resistance matrix $\mathcal{R}$.
@@ 3.2.3 Modified Lü system
In this case, one has $\nabla\times\mathbf{J}\neq 0$ but $\mathbf{J}\cdot(\nabla\times\mathbf{J})=0$, meaning that the Jacobi identity (2.7) is satisfied.
@@ 3.2.4 Qi system
In this case one has $\nabla\times\mathbf{J}\neq 0$ but $\mathbf{J}\cdot(\nabla\times\mathbf{J})=0$, meaning that the Jacobi identity (2.7) is satisfied.
@@ 4.1.2 Lü system
Although the Poisson vector $\mathbf{N}=(N_{x},N_{y},N_{z})$ satisfies the condition $\nabla\cdot(\nabla\times\mathbf{N})=0$ for arbitrary values of the underlying parameters, let us pick $\beta=\gamma$ for the sake of simplicity as this gives $\nabla\times\mathbf{N}=0$ and consequently from (4.5) we get $M=1$ and
@@ 4.2 Jordan-like structures
This opens up certain possibilities; most importantly that of generating new bi-Hamiltonian systems guided by the structure of Jordan homomorphisms [^6] Recall that a map between two rings $\phi:R\rightarrow R^{\prime}$ is a Jordan homomorphism if (a) $\phi(a+b)=\phi(a)+\phi(b)$, (b) $\phi(ab+ba)=\phi(a)\phi(b)+\phi(b)\phi(a)$, for all $a,b\in R$. In the present case, although we do not get a Jordan algebra, we can make use of transformations inspired from Jordan homomorphisms which we will call Jordan-like transformations.. Let us consider an invertible matrix $\mathcal{T}$ under which $\mathcal{J}$ and $\mathcal{R}$ transform as
@@
Note that upon putting $\Delta=1$, one recovers the bi-Hamiltonian system (4.11). The system admits a non-trivial Jacobi last multiplier which may be evaluated straightforwardly and this is therefore not done here. In summary, Jordan-like transformations that map $\mathcal{N}=\mathcal{J}\mathcal{R}+\mathcal{R}\mathcal{J}$ to another Poisson matrix $\mathcal{N}^{\prime}$ can be used to generate new bi-Hamiltonian systems as illustrated by the simple example above.
@@ 5 Conformal Hamiltonian dynamics on Poisson manifolds
$\pounds_{X^{a}_{H}}\Omega=(na)\Omega,$ (5.3)
@@
Note that upon putting $a=0$ one recovers the standard Hamiltonian vector field. The Hamiltonian is not conserved under the flow of $X^{a}_{H}$ for $a\neq 0$ as $X^{a}_{H}(H)=a\Gamma(H)$. Thus, a conformal Hamiltonian vector field has the typical characteristics required for describing a non-conservative time evolution. A key feature of a conformal Hamiltonian vector field is the constancy of its divergence, i.e., although the divergence is non-zero, it is a constant. In what follows, we will demonstrate how some of our previously-considered examples can be described alternatively using conformal Hamiltonian dynamics on Poisson manifolds, i.e., without resorting to resistance matrices.
@@ 5.1 Reduced three-wave-interaction model
The phase trajectories of the system (3.9) are then the integral curves of the conformal Hamiltonian vector field $X^{-2}_{H}$, i.e., with a=2a=-2 in (5.2).
@@ Appendix A Reverse-engineering non-conservative systems
Obviously, the Jacobi identity is satisfied. The Euler equations are obtained from $\dot{\mathbf{x}}=\mathcal{N}(\nabla G)$, where $\mathbf{x}=(L_{x}\quad L_{y}\quad L_{z})^{T}$ and $G$ is the identified with the total kinetic energy $K$. The result is the familiar system of equations